OpenStudy (anonymous):

Someone please help me step by step i've been stuck on this and i want to learn 1. D = p(3 + nr) solve for r

hero (hero):

@milli_lovesyouu begin by dividing both sides by p.

OpenStudy (anonymous):

D/p = (3+nr) ?

hero (hero):

Very good. Now 3 + nr isn't being multiplied by anything anymore so the parentheses are no longer needed. You can remove them.

OpenStudy (anonymous):

okay so then i have to get rid of the n right?

hero (hero):

The next step after removing the parentheses is to subtract 3 from both sides.

OpenStudy (anonymous):

-3d/p = nr ?

hero (hero):

The right side is correct, but the left side is incorrect. You wrote -3 as multiplying d/p, but what you actually what to do is subtract 3 from d/p. Have an idea of how to write that?

OpenStudy (anonymous):

no idea :(

hero (hero):

How do you write six minus 3?

OpenStudy (anonymous):

6 - 3

OpenStudy (anonymous):

3 - d/p ?

hero (hero):

How do you write one-third minus 3?

hero (hero):

Very good. How do you write d/p minus three?

OpenStudy (anonymous):

d/p - 3 = nr?

hero (hero):

Yes, correct. Subtracting 3 from d/p means the same as d/p minus three. You should pay very close attention to the way math phrases are worded. Now the last step is important and I want you to get it right. First I want you to place everything on the left side in parentheses.

OpenStudy (anonymous):

no idea im sorry i tried on paper but it didnt make sense to me

hero (hero):

Type it first in here. I'll explain it later. Just simply place the expression on the left side within a set of parentheses.

OpenStudy (anonymous):

d/p is outside the ( ) right ?

hero (hero):

The entire expression on the left side, put parentheses around it.

OpenStudy (anonymous):

(d/p - 3) = nr then, multiple both sides by n?

hero (hero):

The expression above is correct as written. Notice that n is being multiplied by r. So to undo that, we have to actually perform an inverse operation. DIVISION. Divide both sides by n.

hero (hero):

The final expression is going to look kind of awkward, I know, but no one says the equations have to look pretty.

hero (hero):

You're changing it to something else. When you put the expression in parentheses, YOU CANNOT alter it. It's ONE expression. You divide the entire expression by n.

OpenStudy (anonymous):

okay, so it'll be (d/p - 3) / n = r ?

hero (hero):

Exactly correct.

hero (hero):

And if you're confused about how that looks on paper, it will look like this: \(\dfrac{\dfrac{d}{p} - 3}{n} = r\)

OpenStudy (anonymous):

thats my answer?

hero (hero):

Yes

OpenStudy (anonymous):

these are my answer choices though ?

OpenStudy (anonymous):

@Hero

hero (hero):

It's unfortunate that they have done that, because now they basically want you to re-write it in a different form.

OpenStudy (anonymous):

can you help me? :( please

hero (hero):

No wonder you were confused about how to write the final form.

OpenStudy (anonymous):

yea i was looking at my book and other examples but i couldn't figure it out for nothing

hero (hero):

It will be more difficult to explain based on the manipulation necessary to re-write it in the proper form.

OpenStudy (anonymous):

its okay ill just email my teacher about it i guess

OpenStudy (anonymous):

Thank you so much for all your help!

hero (hero):

Why do you need to email your teacher about it? Have you given up on solving this?

OpenStudy (anonymous):

what other way can i write this?

OpenStudy (anonymous):

i was looking at answer choice a at first

hero (hero):

The next step is to first re-write it this way: \(\dfrac{1}{n}\left(\dfrac{d}{p} - 3\right) = r\) Believe it or not, the expression on the left is equivalent to the expression we got for our initial answer.

hero (hero):

Next, you distribute \(\dfrac{1}{n}\) over \(\dfrac{d}{p} - 3\) to get \(\dfrac{d}{pn} - \dfrac{3}{n}\)

hero (hero):

Afterwards multiply \(\dfrac{3}{n}\) by \(\dfrac{p}{p}\) to get \(\dfrac{3p}{p}\) So the expression on the left now looks like \(\dfrac{d}{pn} - \dfrac{3p}{pn}\

hero (hero):

Notice the denominators are the same so we can combine the expression to get: \(\dfrac{d - 3p}{pn} = r\)

OpenStudy (anonymous):

So, A right?

hero (hero):

A is correct but the important thing to know is how to get A. You can't get it by eyeballing it and you should never try to guess with problems like these.

OpenStudy (anonymous):

yea, i like explanations so i learn because these problems are on my next test :( im in college intermediate Algebra

OpenStudy (anonymous):

Thank you so much!!!!